please help
1.maria travels upstream by boat a distance of 60 miles in 15 hours. the return trip takes 12 hours. find the boats rate in still water and the rate of the current.
2. a truck travels for 4 hours at an average speed of 42 miles per hour. how much longer must the truck travel at an average speed of 55miles per hour so that the average speed of the total trip will be 50 miles per hour?

Answer by mananth(16949)

1.maria travels upstream by boat a distance of 60 miles in 15 hours. the return trip takes 12 hours. find the boats rate in still water and the rate of the current.
let speed of boat in still water be x mph
speed of current = y mph
...
upstream speed will be x-y mph
downstream speed will be x+y mph
...
upstream
60/x-y = 15
60=15(x=y)
60=15x-15y
x-y=4.........................1
...
downstream
60/x+y = 12
60=12(x+y)
60 =12x+12y
x+y=5..........................2
add 1 & 2
2x=9
x= 4.5 mph the speed of boat in still water
y =0.5mph current speed
....
2. a truck travels for 4 hours at an average speed of 42 miles per hour. how much longer must the truck travel at an average speed of 55miles per hour so that the average speed of the total trip will be 50 miles per hour?
let the time it has to travel at 55mph be x hours
4 hours 42 mph . so distance = 168 miles
...
speed = 55 mph time = xhours distance = 55x
...
total distance / total time = speed
55x+168 / (x+4) =50
55x+168=50(x+4)
55x+168=50x+200
55x-50x=200-168
5x=32
x= 6.4 hours
m.ananth@hotmail.ca